Hello!
I wonder what kind of vehicle (small or big) would be best in a head-on car crash with another car. I also wonder where the safest seat is in a head-on car crash, like where could I sit to get as least/lightly damage as I can. I am trying to understand this but I have trouble finding the connections to this with kinetic energy, impulse, momentum, power, and acceleration. I would apprentice if you could help.

Thanks in advance.
Best regards
Zina Gawanson
1
Expert's answer
2020-04-22T10:04:13-0400

"K.E. =\frac12mv^2"

"p=mv"

multiplying 2m on both sides

"K.E\times m\times2=\frac12mv^2\times m\times2"

"2mK.E.=m^2v^2"

substituting the value of p in the above equation we get

"p^2=2mK.E.\\p=\sqrt{2mK.E.}"